# Universal Hitchin moduli spaces

### v1

@@ 3.2. Kähler connection and norm squared of the Higgs field
Let $E_{G}$ be a smooth principal $G$-bundle over $\Sigma$. Let $\curly D$} be the space of connections on $E_{G}$, equipped with the constant complex structure (2.1). Let $\curly{J}$ be the space of complex structures on $\Sigma$ compatible with the given orientation. Consider the space
(3.5) $\curly{X}=\curly{J}\times{\curly D}$
@@
In order to give a more explicit description of $\widehat{\boldsymbol{\omega}}_{\mathbf{J}}$, we consider $h\in\Omega^{0}(E_{G}(G/K))$ a reduction of structure group of $E_{G}$ to $K$, and let $E_{K}$ be the corresponding principal $K$-bundle. From the decomposition $\mathfrak{g}=\mathfrak{k}\oplus i\mathfrak{k}$ one has the identification ${\curly D}\cong{\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k}))$ in (2.7), which induces a biholomorphism
(3.6) $\curly{X}\cong\curly{J}\times{\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k})),$
where the right hand side is endowed with the product complex structure
@@
One can clearly see that $\widehat{\boldsymbol{\omega}}_{\mathbf{J}}$ has a non-trivial dependence on $J\in\curly{J}$, and hence it defines a non-trivial structure of Kähler fibration on $\curly{X}\to\curly{J}$.
The aim of this section is to prove that $(\curly{X}\to\curly{J},\widehat{\boldsymbol{\omega}})_{\mathbf{J}}$ admits a Kähler Ehresmann connection, following Theorem 3.2. For this, we follow a suggestion by N. Hitchin (cf. [Hi1, Section 9]) and define a real $(1,1)$-form $\boldsymbol{\sigma}^{\mathbb{J}}\in\Omega^{1,1}(\curly{X},\mathbb{R})$ by
(3.9) $\boldsymbol{\sigma}^{\mathbb{J}}=i\partial_{\mathbb{J}}\bar{\partial}_{\mathbb{J}}\|\psi\|^{2}_{L^{2}}$
where $\|\psi\|^{2}_{L^{2}}$ denotes the $J$-dependent $L^{2}$-norm of the unitary Higgs field $\psi$, namely
@@
In the next result we calculate an explicit formula for $\boldsymbol{\sigma}^{\mathbb{J}}$.
@@
(3.10) $\begin{split}\boldsymbol{\sigma}^{\mathbb{J}}_{|(J,A,\psi)}((\dot{J}_{1},a_{1},\dot{\psi}_{1}),(\dot{J}_{2},a_{2},\dot{\psi}_{2}))&=\int_{\Sigma}B(a_{1}\wedge(J\dot{\psi}_{2}-\psi(\dot{J}_{2})))\\ &-\int_{\Sigma}B(a_{2}\wedge(J\dot{\psi}_{1}-\psi(\dot{J}_{1})))\\ &-\int_{\Sigma}B\left(\left(J\dot{\psi}_{1}-\tfrac{1}{2}\psi(\dot{J}_{1})\right)\wedge\psi(\dot{J}_{2})\right)\\ &+\int_{\Sigma}B\left(\left(J\dot{\psi}_{2}-\tfrac{1}{2}\psi(\dot{J}_{2})\right)\wedge\psi(\dot{J}_{1})\right)\end{split}$
@@
Given $C\in\operatorname{End}T\Sigma$ and $\psi_{1},\psi_{2}\in\Omega^{1}(X,E_{K}(\mathfrak{k}))$, the following pointwise identity holds:
@@
By direct calculation at a point $x\in\Sigma$, if $C=v\otimes\gamma\in T_{x}\Sigma\otimes T^{*}_{x}\Sigma$, by dimensional reasons we have
@@
$=\gamma\wedge(B(\psi_{1}(v),\psi_{2})-B(\psi_{1},\psi_{2}(v)))=B(\psi_{1}(v)\gamma\wedge\psi_{2})+B(\psi_{1}\wedge\psi_{2}(v)\gamma).$
Our next result proves the existence of a natural Kähler Ehresmann connection in our Kähler fibration $\curly{X}\to\curly{J}$, with coupling form $\boldsymbol{\sigma}^{\mathbb{J}}$ and curvature given (essentially) by the unitary Higgs field $\psi$.
@@
where $\sigma$ is defined by (3.9). More explicitly,
@@
Furthermore, $H^{\mathbb{J}}$ is preserved by $\mathbb{J}$ and the curvature $F_{\mathbb{J}}:=F_{\Gamma^{\mathbb{J}}}\in\Omega^{2}(\curly{X},V\curly{X})$ of $\Gamma$ is of type $(1,1)$ and given explicitly by
(3.12) $\begin{split}(F_{\mathbb{J}})_{|(J,A,\psi)}(v_{1},v_{2})&=(0,\psi([\dot{J}_{2},\dot{J}_{1}]),0).\end{split}$
@@
(3.13) $\begin{split}\boldsymbol{\sigma}^{\mathbb{J}}((0,a_{1},\dot{\psi}_{1}),(0,a_{2},\dot{\psi}_{2}))&=\int_{\Sigma}B(a_{1}\wedge J\dot{\psi}_{2})+\int_{\Sigma}B(J^{2}\dot{\psi}_{1}\wedge Ja_{2})\\ &=\widehat{\boldsymbol{\omega}}_{\mathbf{J}}((0,a_{1},\dot{\psi}_{1}),(0,a_{2},\dot{\psi}_{2})).\end{split}$
where we ommit the evaluation at the point $(J,A,\psi)$ for simplicity in the notation. Contracting now $\boldsymbol{\sigma}^{\mathbb{J}}$ with a vertical vector field
$\begin{split}\boldsymbol{\sigma}^{\mathbb{J}}((0,a_{1},\dot{\psi}_{1}),(\dot{J}_{2},a_{2},\dot{\psi}_{2}))&=\int_{\Sigma}B(a_{1}\wedge(J\dot{\psi}_{2}-\psi(\dot{J}_{2})))\\ &-\int_{\Sigma}B\left(J\dot{\psi}_{1}\wedge(\psi(\dot{J}_{2})-a_{2})\right)\end{split}$
and therefore $(\dot{J}_{2},a_{2},\dot{\psi}_{2}))\in H^{\mathbb{J}}$ if and only if
@@
$\begin{split}\boldsymbol{\sigma}^{\mathbb{J}}(v_{1},v_{2})&=-\int_{\Sigma}B\left(\left(\psi(\dot{J}_{1})-\tfrac{1}{2}\psi(\dot{J}_{1})\right)\wedge\psi(\dot{J}_{2})\right)+\int_{\Sigma}B\left(\left(\psi(\dot{J}_{2})-\tfrac{1}{2}\psi(\dot{J}_{2})\right)\wedge\psi(\dot{J}_{1})\right)\\ &=-\int_{\Sigma}B(\psi(\dot{J}_{1}))\wedge\psi(\dot{J}_{2}))\end{split}$
@@
$d(\boldsymbol{\sigma}^{\mathbb{J}}(v_{1},v_{2}))((0,a,\dot{\psi}))=-\int_{\Sigma}B(\dot{\psi}(\dot{J}_{1}))\wedge\psi(\dot{J}_{2}))-\int_{\Sigma}B(\psi(\dot{J}_{1}))\wedge\dot{\psi}(\dot{J}_{2})).$
On the other hand, applying Lemma 3.5,
$\widehat{\boldsymbol{\omega}}_{\mathbf{J}}((0,(J\psi)([\dot{J}_{1},\dot{J}_{2}],0),(0,a,\dot{\psi}))$ $=\int_{\Sigma}B((J\psi)([\dot{J}_{1},\dot{J}_{2}])\wedge J\dot{\psi})$
@@
$=d(\boldsymbol{\sigma}^{\mathbb{J}}(v_{1},v_{2}))((0,a,\dot{\psi})),$
which proves the last part of the statement, by application of Proposition 3.3. ∎
@@
(3.14) $\begin{split}\boldsymbol{\sigma}^{\mathbb{J}}(v_{1},v_{2})&=\int_{\Sigma}B((a_{1}-\psi(\dot{J}_{1}))\wedge J(\dot{\psi}_{2}-(J\psi)(\dot{J}_{2})))\\ &-\int_{\Sigma}B((\dot{\psi}_{1}-(J\psi)(\dot{J}_{1}))\wedge J(a_{2}-\psi(\dot{J}_{2})))\\ &-\int_{\Sigma}B(\psi(\dot{J}_{1}))\wedge\psi(\dot{J}_{2})),\\ \mathbf{g}_{\mathbb{J}}(v,v)&=\int_{\Sigma}B((a-\psi(\dot{J}))\wedge J(a-\psi(\dot{J})))\\ &+\int_{\Sigma}B((\dot{\psi}-(J\psi)(\dot{J}))\wedge J(\dot{\psi}-(J\psi)(\dot{J}))\\ &-\int_{\Sigma}B(\psi(\dot{J}))\wedge J(\psi(\dot{J}))).\end{split}$
@@
(3.15) $\begin{split}\mathbf{g}_{\mathbb{J}}(v,v)=-\int_{\Sigma}B\left(\psi(\dot{J})\wedge J(\psi(\dot{J}))\right).\end{split}$
@@
3.3. Minimal coupling and Hamiltonian action
Let $(\curly{X}\to\curly{J},\widehat{\boldsymbol{\omega}}_{\mathbf{J}})$ be the Kähler fibration defined in Section 3.2. In this section we undertake minimal couplings for the connection $\Gamma^{\mathbb{J}}$ constructed in Proposition 3.6. We will see that the corresponding (pre)symplectic structures on $\curly{X}$ admit a Hamiltonian action by a suitable extended gauge group $\widetilde{{\curly K}}$, naturally associated to a choice of symplectic structure $\omega$ on $\Sigma$. This will allow us, in Section 4, to construct pseudo-Kähler structures on the universal moduli space of solutions of the harmonicity equations (2.14) over Teichmuller space.
We start by introducing the Kähler structure we shall consider, up to sign, in the base of of the fibration $\curly{X}\to\curly{J}$. Following Donaldson [Do2] and Fujiki [F], we fix a symplectic (volume) form $\omega$ on $\Sigma$ compatible with the given orientation. The space $\curly{J}$ of complex structures $J$ on $\Sigma$ compatible with the orientation (and hence with $\omega$) is an infinite dimensional Kähler manifold, with complex structure $\mathbb{J}{J}\colon T_{J}\curly{J}\to T_{J}\curly{J}$ and Kähler form $\boldsymbol{\omega}{J}$ given, respectively, by
@@ 3.3. Minimal coupling and Hamiltonian action
for $\dot{J}_{1},\dot{J}_{2}\in T_{J}\curly{J}$. Note that, by dimensional reasons, any $\dot{J}\in T_{J}\curly{J}$, regarded as an endomorphism $\dot{J}\colon T\Sigma\to T\Sigma$, is symmetric with respect to the induced metric $g=\omega(\cdot,J\cdot)$
Let $(\curly{X}\to\curly{J},\widehat{\boldsymbol{\omega}}_{\mathbf{J}})$ be the Kähler fibration defined in Section 3.2. Given $\alpha>0$ a real coupling constant and $\varepsilon\in\{-1,1\}$, the family of minimal coupling symplectic structures of our interest is defined by
@@
Applying Corollary 3.7, $\boldsymbol{\omega}_{\alpha}^{-1}$ is negative-definite along the horizontal subspace $H^{\mathbb{J}}\subset T\curly{X}$. The analysis for the case $\varepsilon=1$ is much more subtle, as we can see from the next result.
@@
Given a horizontal vector field $v\in H^{\mathbb{J}}$ at $(J,A,\psi)$, covering $\dot{J}\in T_{J}\curly{J}$, one has
@@
    For $\varepsilon=-1$, $\mathbf{g}_{\alpha}^{\varepsilon}$ is negative definite along $H^{\mathbb{J}}$. In particular, $\boldsymbol{\omega}_{\alpha}^{-1}$ is a non-degenerate symplectic structure.
@@
    For $\varepsilon=1$ and $\psi\neq 0$, $\mathbf{g}_{\alpha}^{\varepsilon}$ changes signature along the line $(J,A,\lambda\psi)\in\curly{X}$, for $\lambda\in\mathbb{R}$.
@@
The case $\varepsilon=-1$ is a direct consequence of Corollary 3.7. For the case $\varepsilon=1$, let $(J,A,\psi)\in\curly{X}$. Then, for any $\dot{J}\in T_{J}\curly{J}$, we have a pointwise equalities
@@
The result follows now from the previous cases.
@@
where $p$ maps each $g\in\widetilde{{\curly K}}$ into the Hamiltonian symplectomorphism $p(g)\in\curly{H}$ that it covers, and so its kernel $\curly K$} is the gauge group of $E_{K}$, that is, the normal subgroup of $\widetilde{{\curly K}}$ consisting of automorphisms of $E_{K}$ covering the identity map on $\Sigma$.
Similarly as in [AGG, Section 2.2], the group $\widetilde{{\curly K}}$ acts on $(\curly{X},\mathbb{J},\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon})$ preserving $\mathbb{J}$ and $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$, and covering the $\curly{H}$-action on $\curly{J}$ by push-forward. This last action preserves the Kähler structure on $\curly{J}$, and furthermore is Hamiltonian, with equivariant moment map $\mu\colon\curly{J}\to(\operatorname{Lie}\curly{H})^{*}$ given by
@@
for $f\in C^{\infty}_{0}(\Sigma)$, identified with an element $\eta_{f}$ in $\operatorname{Lie}\curly{H}$, and $S_{J}$ is the scalar curvature of the metric $g=\omega(,J)$ defined by $\omega$ and $J$. Here, $C^{\infty}_{0}(\Sigma)$ denotes the space of smooth functions on $\Sigma$ such that $\int_{\Sigma}f\omega=0$. This fact was shown by Donaldson [Do2] and Fujiki [F] independently for symplectic manifolds of arbitrary dimensions, and it seems that was known to Quillen in the case of surfaces.
@@
where $A\zeta\in\Omega^{0}(X,E_{K}(\mathfrak{k}))$ denotes the vertical part of the $K$-equivariant vector field $\zeta$ on the total space of $E_{K}$, with respect to the connection $A$.
@@
Since $\sigma=\tfrac{1}{2}dd^{c}_{\mathbb{J}}\nu$, with $\nu(J,A,\psi):=\|\psi\|^{2}_{L^{2}}$ (see Lemma 3.4), and $d^{c}_{\mathbb{J}}\nu$ is preserved by the $\widetilde{{\curly K}}-$action, there exists an equivariant moment map given by
@@
where $\zeta\cdot(J,A,\psi)$ is the the infinitesimal action of $\zeta$ on $(J,A,\psi)$. Using the complex connection $D=A+i\psi$ it follows that

### v2

@@ 3.2. The Kähler connection and the squared norm of the Higgs field
Let $E_{G}$ be a smooth principal $G$-bundle over $\Sigma$. Let $\curly D$} be the space of connections on $E_{G}$, equipped with the constant complex structure (2.1), and $\curly{J}$ the space of complex structures on $\Sigma$ that are compatible with the given orientation. Consider the space
(3.5) $\curly{X}=\curly{J}\times{\curly D},$
@@
In order to give a more explicit description of $\widehat{\boldsymbol{\omega}}_{\mathbf{J}}$, we consider a reduction $h\in\Omega^{0}(E_{G}(G/K))$ of the structure group of $E_{G}$ to $K$, and let $E_{K}$ be the corresponding principal $K$-bundle. From the decomposition $\mathfrak{g}=\mathfrak{k}\oplus i\mathfrak{k}$ one has the identification ${\curly D}\cong{\curly A}\times\Omega^{1}(\Sigma,E_{K}(\mathfrak{k}))$ in (2.6), which induces a biholomorphism
(3.6) $\curly{X}\cong\curly{J}\times{\curly A}\times\Omega^{1}(\Sigma,E_{K}(\mathfrak{k})),$
where the right-hand side is endowed with the product complex structure
@@
It is clear that $\widehat{\boldsymbol{\omega}}_{\mathbf{J}}$ has a non-trivial dependence on $J\in\curly{J}$, and hence defines a structure of a non-trivial Kähler fibration on $\curly{X}\to\curly{J}$.
The aim of this section is to prove that $(\curly{X}\to\curly{J},\widehat{\boldsymbol{\omega}}_{\mathbf{J}})$ admits a Kähler Ehresmann connection, following Theorem 3.2. For this, we follow a suggestion by N. Hitchin (cf. [Hi1, Section 9]) and define a real $(1,1)$-form $\boldsymbol{\sigma}^{\mathbb{J}}\in\Omega^{1,1}(\curly{X},\mathbb{R})$ by
(3.9) $\boldsymbol{\sigma}^{\mathbb{J}}=i\partial_{\mathbb{J}}\bar{\partial}_{\mathbb{J}}\|\psi\|^{2}_{L^{2}},$
where $\|\psi\|^{2}_{L^{2}}$ denotes the $J$-dependent $L^{2}$-norm of the ‘unitary Higgs field’ $\psi$, namely
@@
In the next result, we derive an explicit formula for $\boldsymbol{\sigma}^{\mathbb{J}}$.
@@
(3.10) $\begin{split}\boldsymbol{\sigma}^{\mathbb{J}}_{|(J,A,\psi)}((\dot{J}_{1},a_{1},\dot{\psi}_{1}),(\dot{J}_{2},a_{2},\dot{\psi}_{2}))&=\int_{\Sigma}B(a_{1}\wedge(J\dot{\psi}_{2}-\psi(\dot{J}_{2})))\\ &-\int_{\Sigma}B(a_{2}\wedge(J\dot{\psi}_{1}-\psi(\dot{J}_{1})))\\ &-\int_{\Sigma}B((J\dot{\psi}_{1}-\tfrac{1}{2}\psi(\dot{J}_{1}))\wedge\psi(\dot{J}_{2}))\\ &+\int_{\Sigma}B((J\dot{\psi}_{2}-\tfrac{1}{2}\psi(\dot{J}_{2}))\wedge\psi(\dot{J}_{1}))\end{split}$
@@
For $C\in\operatorname{End}T\Sigma$ and $\psi_{1},\psi_{2}\in\Omega^{1}(\Sigma,E_{K}(\mathfrak{k}))$, the following pointwise identity holds:
@@
By direct calculation at a point $x\in\Sigma$, if $C=v\otimes\gamma\in T_{x}\Sigma\otimes T^{*}_{x}\Sigma$, we see by dimensional reasons that
@@
$=\gamma\wedge(B(\psi_{1}(v),\psi_{2})-B(\psi_{1},\psi_{2}(v)))=B(\psi_{1}(v)\gamma\wedge\psi_{2})+B(\psi_{1}\wedge\psi_{2}(v)\gamma).\qed$
Our next result proves the existence of a natural Kähler Ehresmann connection on the Kähler fibration $\curly{X}\to\curly{J}$, whose coupling form is $\boldsymbol{\sigma}^{\mathbb{J}}$ and whose curvature is essentially given by the ‘unitary Higgs field’ $\psi$.
@@
where $\boldsymbol{\sigma}^{\mathbb{J}}$ is defined by (3.9). More explicitly,
@@
Furthermore, $H^{\mathbb{J}}$ is preserved by $\mathbb{J}$ and the curvature $F_{\mathbb{J}}:=F_{\Gamma^{\mathbb{J}}}\in\Omega^{2}(\curly{X},V\curly{X})$ of $\Gamma^{\mathbb{J}}$ is of type $(1,1)$, and given explicitly by
(3.12) $(F_{\mathbb{J}})_{|(J,A,\psi)}(v_{1},v_{2})=(0,\psi([\dot{J}_{2},\dot{J}_{1}]),0),$
@@
(3.13) $\begin{split}\boldsymbol{\sigma}^{\mathbb{J}}((0,a_{1},\dot{\psi}_{1}),(0,a_{2},\dot{\psi}_{2}))&=\int_{\Sigma}B(a_{1}\wedge J\dot{\psi}_{2})+\int_{\Sigma}B(J^{2}\dot{\psi}_{1}\wedge Ja_{2})\\ &=\widehat{\boldsymbol{\omega}}_{\mathbf{J}}((0,a_{1},\dot{\psi}_{1}),(0,a_{2},\dot{\psi}_{2})),\end{split}$
where we omit the evaluation at the point $(J,A,\psi)$ for simplicity in the notation. Contracting now $\boldsymbol{\sigma}^{\mathbb{J}}$ with a vertical vector field
$\begin{split}\boldsymbol{\sigma}^{\mathbb{J}}((0,a_{1},\dot{\psi}_{1}),(\dot{J}_{2},a_{2},\dot{\psi}_{2}))&=\int_{\Sigma}B(a_{1}\wedge(J\dot{\psi}_{2}-\psi(\dot{J}_{2})))-\int_{\Sigma}B(J\dot{\psi}_{1}\wedge(\psi(\dot{J}_{2})-a_{2})),\end{split}$
and therefore $(\dot{J}_{2},a_{2},\dot{\psi}_{2})\in H^{\mathbb{J}}$ if and only if
@@
$\begin{split}\boldsymbol{\sigma}^{\mathbb{J}}(v_{1},v_{2})&=-\int_{\Sigma}B((\psi(\dot{J}_{1})-\tfrac{1}{2}\psi(\dot{J}_{1}))\wedge\psi(\dot{J}_{2}))+\int_{\Sigma}B((\psi(\dot{J}_{2})-\tfrac{1}{2}\psi(\dot{J}_{2}))\wedge\psi(\dot{J}_{1}))\\ &=-\int_{\Sigma}B(\psi(\dot{J}_{1})\wedge\psi(\dot{J}_{2})),\end{split}$
@@
$d(\boldsymbol{\sigma}^{\mathbb{J}}(v_{1},v_{2}))((0,a,\dot{\psi}))=-\int_{\Sigma}B(\dot{\psi}(\dot{J}_{1})\wedge\psi(\dot{J}_{2}))-\int_{\Sigma}B(\psi(\dot{J}_{1})\wedge\dot{\psi}(\dot{J}_{2})).$
On the other hand, applying Lemma 3.5, we obtain
$\widehat{\boldsymbol{\omega}}_{\mathbf{J}}((0,(J\psi)([\dot{J}_{1},\dot{J}_{2}]),0),(0,a,\dot{\psi}))$ $=\int_{\Sigma}B((J\psi)([\dot{J}_{1},\dot{J}_{2}])\wedge J\dot{\psi})$
@@
$=d(\boldsymbol{\sigma}^{\mathbb{J}}(v_{1},v_{2}))(0,a,\dot{\psi}),$
which proves the last part of the statement, by Proposition 3.3. ∎
@@
(3.14) $\begin{split}\boldsymbol{\sigma}^{\mathbb{J}}(v_{1},v_{2})&=\int_{\Sigma}B((a_{1}-\psi(\dot{J}_{1}))\wedge J(\dot{\psi}_{2}-(J\psi)(\dot{J}_{2})))\\ &-\int_{\Sigma}B((\dot{\psi}_{1}-(J\psi)(\dot{J}_{1}))\wedge J(a_{2}-\psi(\dot{J}_{2})))\\ &-\int_{\Sigma}B(\psi(\dot{J}_{1})\wedge\psi(\dot{J}_{2})),\\ \mathbf{g}_{\mathbb{J}}(v,v)&=\int_{\Sigma}B((a-\psi(\dot{J}))\wedge J(a-\psi(\dot{J})))\\ &+\int_{\Sigma}B((\dot{\psi}-(J\psi)(\dot{J}))\wedge J(\dot{\psi}-(J\psi)(\dot{J})))\\ &-\int_{\Sigma}B(\psi(\dot{J})\wedge J(\psi(\dot{J}))).\end{split}$
@@
(3.15) $\begin{split}\mathbf{g}_{\mathbb{J}}(v,v)=-\int_{\Sigma}B(\psi(\dot{J})\wedge J(\psi(\dot{J}))).\end{split}$
@@
3.3. Minimal coupling and the Hamiltonian action
Let $(\curly{X}\to\curly{J},\widehat{\boldsymbol{\omega}}_{\mathbf{J}})$ be the Kähler fibration defined in Section 3.2. In this section we apply the minimal-coupling construction to the connection $\Gamma^{\mathbb{J}}$ introduced in Proposition 3.6. We will see that the corresponding (pre)symplectic structures on $\curly{X}$ admit a Hamiltonian action by a suitable extended gauge group $\widetilde{{\curly K}}$, naturally associated to a choice of symplectic structure $\omega$ on $\Sigma$. This will allow us, in Section 4, to construct pseudo-Kähler structures on the universal moduli space of solutions of the harmonicity equations (2.11) over the Teichmüller space.
We start by introducing the Kähler structure that we will consider, up to sign, on the base of the fibration $\curly{X}\to\curly{J}$. Following Donaldson [Do2] and Fujiki [Fu], we fix a symplectic (volume) form $\omega$ on $\Sigma$ compatible with the given orientation. The space $\curly{J}$ of complex structures $J$ on $\Sigma$ compatible with the orientation (and hence with $\omega$) is an infinite-dimensional Kähler manifold, with complex structure $\mathbb{J}{J}\colon T_{J}\curly{J}\to T_{J}\curly{J}$ and Kähler form $\boldsymbol{\omega}{J}$ given, respectively, by
@@ 3.3. Minimal coupling and the Hamiltonian action
for $\dot{J}_{1},\dot{J}_{2}\in T_{J}\curly{J}$. Note that, for dimensional reasons, any $\dot{J}\in T_{J}\curly{J}$, regarded as an endomorphism $\dot{J}\colon T\Sigma\to T\Sigma$, is symmetric with respect to the induced metric $g=\omega(\cdot,J\cdot)$.
Let $(\curly{X}\to\curly{J},\widehat{\boldsymbol{\omega}}_{\mathbf{J}})$ be the Kähler fibration defined in Section 3.2. Given a real ‘coupling constant’ $\alpha>0$ and $\varepsilon\in\{-1,1\}$, the family of minimal coupling symplectic structures of our interest is defined by
@@
Applying Corollary 3.7, $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,-1}$ is negative-definite along the horizontal subspace $H^{\mathbb{J}}\subset T\curly{X}$. The analysis for the case $\varepsilon=1$ is much more subtle, as we can see from the next result.
@@
Let $v\in H^{\mathbb{J}}$ be a horizontal vector field at $(J,A,\psi)$, covering $\dot{J}\in T_{J}\curly{J}$. Then
@@
    If $\varepsilon=-1$, $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$ is negative definite along $H^{\mathbb{J}}$. In particular, $\boldsymbol{\omega}_{\alpha}^{-1}$ is a non-degenerate symplectic structure.
@@
    If $\varepsilon=1$ and $\psi\neq 0$, $\mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}$ changes signature along the line $(J,A,\lambda\psi)\in\curly{X}$, for $\lambda\in\mathbb{R}$.
@@
The case $\varepsilon=-1$ is a direct consequence of Corollary 3.7. For the case $\varepsilon=1$, let $(J,A,\psi)\in\curly{X}$. Then, for any $\dot{J}\in T_{J}\curly{J}$, we have pointwise identities
@@
The result follows now from the previous cases. ∎
@@
where $p$ maps each $g\in\widetilde{{\curly K}}$ to the Hamiltonian symplectomorphism $p(g)\in\curly{H}$ that it covers, and so its kernel $\curly K$} is the gauge group of $E_{K}$, that is, the normal subgroup of $\widetilde{{\curly K}}$ consisting of automorphisms of $E_{K}$ covering the identity map on $\Sigma$.
As in [AGG1, Section 2.2], the group $\widetilde{{\curly K}}$ acts on $(\curly{X},\mathbb{J},\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon})$, preserving both $\mathbb{J}$ and $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$, and covering the $\curly{H}$-action on $\curly{J}$ given by push-forward. This last action preserves the Kähler structure on $\curly{J}$ and is moreover Hamiltonian, with equivariant moment map $\mu\colon\curly{J}\to(\operatorname{Lie}\curly{H})^{*}$ given by
@@
where $f\in C^{\infty}_{0}(\Sigma)$ is identified with an element $\eta_{f}$ in $\operatorname{Lie}\curly{H}$, and $S_{J}$ is the scalar curvature of the metric $g=\omega(,J)$ determined by $\omega$ and $J$. Here, $C^{\infty}_{0}(\Sigma)$ denotes the space of smooth functions on $\Sigma$ such that $\int_{\Sigma}f\omega=0$. This fact was established by Donaldson [Do2] and Fujiki [Fu] independently for symplectic manifolds of arbitrary dimension, and it appears to have been known to Quillen in the case of surfaces.
@@
where $A\zeta\in\Omega^{0}(\Sigma,E_{K}(\mathfrak{k}))$ denotes the vertical part of the $K$-equivariant vector field $\zeta$ on the total space of $E_{K}$, with respect to the connection $A$.
@@
Since $\sigma=\tfrac{1}{2}dd^{c}_{\mathbb{J}}\nu$, with $\nu(J,A,\psi):=\|\psi\|^{2}_{L^{2}}$ (see Lemma 3.4), and because $d^{c}_{\mathbb{J}}\nu$ is preserved by the $\widetilde{{\curly K}}-$action, there exists an equivariant moment map given by
@@
where $\zeta\cdot(J,A,\psi)$ denotes the infinitesimal action of $\zeta$ on $(J,A,\psi)$. Using the complex connection $D=A+i\psi$, it follows that